OpenStudy (anonymous):

find the vertex,x and y intecept f(x)=3x^2+12x+14

OpenStudy (anonymous):

To find the vertex, you will need to use the vertex formula which is x = -b/ 2a. For this problem, the vertex is -1. To find x-intercepts, you will need to let y=0 and solve for x and let x=0 and solve for y to find the y-intercept.

OpenStudy (anonymous):

x = -b/2a = -12/(2*3) = -2 Refer to the attachment.